How do you work this problem: 8a-3-5a=15
Answer by laurien(6)

Hello-
you need to combine like terms first. The terms 8a and -5a share the a so 8a-5a=3a
3a-3 = 15. You need to add 3 to the -3 to get zero before you can do the same thing on the other side of the = sign Now 15 + 3 = 18 so 3a = 18. The inverse of multiplication 3 * a is division so you divide 3a by 3 to get 1a alone...then divide 18 by 3 to get the solution...a = 6. You can plug 6 in for a to be sure it is correct..8(6) - 3 - (5*6) = 15 so 48-3-30 = 15!
Think Math!....:)